# All or None: Identifiable Linear Properties of Next-token Predictors in Language Modeling

### v1

@@ B.3 Extended linear equivalence
Notice that ${\bm{\mathrm{M}}}^{+}{\bm{\mathrm{M}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}$ and ${\bm{\mathrm{N}}}^{+}{\bm{\mathrm{N}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}$, which follows by the fact that $\mathrm{ker}({\bm{\mathrm{M}}})^{\perp}=\tilde{}\mathcal{M}$ and $\mathrm{ker}({\bm{\mathrm{N}}})^{\perp}=\tilde{}\mathcal{N}$ [60, Page 211]. Using this we have ${\bm{\mathrm{M}}}^{+}{\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}$, and similarly ${\bm{\mathrm{N}}}^{+}{\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}$. Therefore, on the right-hand side only the projectors ${\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}$ and ${\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}$ remain, whereas we have to set $\tilde{{\bm{\mathrm{M}}}}={\bm{\mathrm{M}}}^{+}$ and $\tilde{{\bm{\mathrm{N}}}}={\bm{\mathrm{N}}}^{+}$. Both $\tilde{\bm{\mathrm{M}}}$ and $\tilde{\bm{\mathrm{N}}}$ have range $k$, as a consequence of being pseudo-inverses. Therefore, we get $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})\sim_{EL}({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$.
@@ B.4 Proof of Proposition 4
Starting from expression of ${\bm{\mathrm{f}}}$ and ${\bm{\mathrm{g}}}_{0}$ given by the $\sim_{EL}$ equivalence relation (Definition 3):
@@
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)$ $={\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)$
$={\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)$
$={\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)\,,$
where we used the idempotency of projectors, i.e., that ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}^{2}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}$ and ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}^{2}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}$. ∎
B.5 Proof of Theorem 5
@@
For two matrices ${\bm{\mathrm{A}}}\in\mathbb{R}^{m\times k}$ and ${\bm{\mathrm{B}}}\in\mathbb{R}^{k\times n}$, from the rank-nullity theorem [60, page 62], it follows that:
@@
    Starting from log-equality of probabilities, we adopt a pivoting strategy to get rid of normalizing constants [^5] A similar “pivoting” strategy is the starting point of several identifiability proofs in nonlinear ICA with auxiliary variables, e.g., [43, 24, 20, 3].;
@@
Step 2: Obtaining the relation between embeddings ${\bm{\mathrm{f}}}$ and $\tilde{\bm{\mathrm{f}}}$. We will now show that ${\bm{\mathrm{M}}}:=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}({\bm{\mathrm{G}}}^{\top})^{+}\tilde{{\bm{\mathrm{G}}}}^{\top}({\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}})$ and $\tilde{{\bm{\mathrm{M}}}}:=({\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}})^{+}(\tilde{\bm{\mathrm{G}}}^{\top})^{+}{{\bm{\mathrm{G}}}}^{\top}{\bm{\mathrm{P}}}_{{\mathcal{G}}}{\bm{\mathrm{P}}}_{{\mathcal{F}}}$ satisfies 9, showing that ${\bm{\mathrm{f}}}$ and $\tilde{{\bm{\mathrm{f}}}}$ are linearly related. For this, consider $q$ elements, or tokens, $y\in\mathcal{A}$ such that the matrix
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{G}}})}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $=({\bm{\mathrm{G}}}^{\top})^{+}\tilde{{\bm{\mathrm{G}}}}^{\top}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$ (From 30, $({\bm{\mathrm{G}}}^{\top})^{+}{\bm{\mathrm{G}}}^{\top}={\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{G}}})}$)
@@
    We show that $m\leq k$ from the definition of ${\bm{\mathrm{M}}}$ in Equation 98;
@@
    We show that necessarily $m\geq k$ and $k^{\prime}\geq k$ from Equation 97;
@@
    We show that $k^{\prime}\leq k$ from Equation 99.
(I) By equation (98), we have that ${\bm{\mathrm{M}}}={\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{M}}}$, since by 98, in ${\bm{\mathrm{M}}}$ we have on the left ${\bm{\mathrm{P}}}_{\mathcal{M}}({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}$ by Lemma 2 (ii). Taking the rank of ${\bm{\mathrm{M}}}$ and using the fact that $\mathrm{rank}({\bm{\mathrm{A}}}{\bm{\mathrm{B}}})\leq\min\big{(}\mathrm{rank}({\bm{\mathrm{A}}}),\mathrm{rank}({\bm{\mathrm{B}}})\big{)}$ [60], we obtain:
@@
Together with 101, it shows that $k\leq m\leq k$, so it must be that $m=k$ and so $\mathrm{rank}({\bm{\mathrm{M}}})=\mathrm{dim}(\mathcal{M})$.
@@
which, together with 106, it holds only as long as $k^{\prime}\leq k$. This shows that $k\leq k^{\prime}\leq k$, so it must be that $k^{\prime}=k$.
@@
Step 4. Obtaining the relation between unembeddings ${\bm{\mathrm{g}}}$ and $\tilde{\bm{\mathrm{g}}}$. We will now show that ${\bm{\mathrm{N}}}:=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}({\bm{\mathrm{M}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}$, using the matrix ${\bm{\mathrm{M}}}$ in 98 from Step 2 satisfies Equation 10.
Similar to Step 2, we take $\ell$ points, or sequences, ${\bm{\mathrm{x}}}_{i}\in\mathsf{Seq}(\mathcal{A})$ such that ${\bm{\mathrm{F}}}$ and $\tilde{{\bm{\mathrm{F}}}}$ in Equation 102 span $\mathcal{F}$ and $\tilde{\mathcal{F}}$, respectively. We then have:
@@
${\bm{\mathrm{F}}}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}}{\bm{\mathrm{g}}}_{0}(y)$ $=\tilde{{\bm{\mathrm{F}}}}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (Using orthogonal projectors 93)
@@
where in the last line we substituted the expression for ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{F}}}$ given by 98. Thus, restarting from (110), and reworking the expression we get:
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}}{\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{0}(y)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (From 30 we get $(\tilde{{\bm{\mathrm{F}}}}^{\top})^{+}(\tilde{{\bm{\mathrm{F}}}}^{\top})={\bm{\mathrm{P}}}_{\mathrm{Im}(\tilde{{\bm{\mathrm{F}}}})}={\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}$)
@@
where we used in the last line that ${\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}{\bm{\mathrm{M}}}^{\top}={\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}{\bm{\mathrm{M}}}^{\top}={\bm{\mathrm{M}}}^{\top}$, and ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{M}}}^{\top}={\bm{\mathrm{M}}}^{\top}$ follows by the definition of ${\bm{\mathrm{M}}}$ 98, containing on the right $({\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}})$. Recall that by Lemma 2 (iii), it holds that
@@
Using this in 111, we obtain
@@
where in the last line we use $({\bm{\mathrm{M}}}^{\top})^{+}{\bm{\mathrm{M}}}^{\top}={\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{M}}})}={\bm{\mathrm{P}}}_{\mathcal{M}}$, where the first equality follows by 30 and the second one is given by $\mathrm{Im}({\bm{\mathrm{M}}})=\mathcal{M}$ from 109. We now use that ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}={\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}$ to obtain:
@@
    We use 116 to obtain that $\mathrm{rank}({\bm{\mathrm{N}}})\geq k$.
@@
(II) Next, consider $q$ elements of $\mathcal{A}$ such that the matrices ${\bm{\mathrm{G}}}$ and $\tilde{{\bm{\mathrm{G}}}}$ in Equation 91 have rank equal to $\mathrm{dim}(G)$ and $\mathrm{dim}(\tilde{\mathcal{G}})$, respectively. From Equation 116, it holds:

### v2

@@ B.3 Extended linear equivalence
Notice that ${\bm{\mathrm{M}}}^{+}{\bm{\mathrm{M}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}$ and ${\bm{\mathrm{N}}}^{+}{\bm{\mathrm{N}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}$, which follows by the fact that $\mathrm{ker}({\bm{\mathrm{M}}})^{\perp}=\tilde{}\mathcal{M}$ and $\mathrm{ker}({\bm{\mathrm{N}}})^{\perp}=\tilde{}\mathcal{N}$ [Axler, 2015, Page 211]. Using this we have ${\bm{\mathrm{M}}}^{+}{\bm{\mathrm{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}$, and similarly ${\bm{\mathrm{N}}}^{+}{\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}={\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}$. Therefore, on the right-hand side only the projectors ${\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}$ and ${\bm{\mathrm{P}}}_{\tilde{\mathcal{N}}}$ remain, whereas we have to set $\tilde{{\bm{\mathrm{M}}}}={\bm{\mathrm{M}}}^{+}$ and $\tilde{{\bm{\mathrm{N}}}}={\bm{\mathrm{N}}}^{+}$. Both $\tilde{\bm{\mathrm{M}}}$ and $\tilde{\bm{\mathrm{N}}}$ have range $k$, as a consequence of being pseudo-inverses. Therefore, we get $(\tilde{{\bm{\mathrm{f}}}},\tilde{{\bm{\mathrm{g}}}})\sim_{EL}({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$.
@@ B.4 Proof of Proposition 4
Starting from Lemma 2 (vi), we have that:
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)={\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)\,.$ (80)
Considering the expression of ${\bm{\mathrm{f}}}$ and ${\bm{\mathrm{g}}}_{0}$ given by the $\sim_{EL}$ equivalence relation (Definition 3):
@@
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{N}}{\bm{\mathrm{g}}}_{0}(y)$ $=\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{N}}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)$
$=\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)$
$=\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)\,,$
where we used the idempotency of projectors, i.e., that ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}^{2}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}$ and ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}^{2}={\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}$. To prove the claim, it is sufficient to use Lemma 2 (vi) again, obtaining:
${\bm{\mathrm{f}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{g}}}_{0}(y)$ $=\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})^{\top}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{N}}\tilde{\bm{\mathrm{g}}}_{0}(y)=\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})^{\top}\tilde{\bm{\mathrm{g}}}_{0}(y)\,.$ (83)
B.5 Counterexample when diversity condition does not hold
We detail here a counter-example to linear identifiability (Corollary 6) when the diversity condition does not hold. Let ${\bm{\mathrm{f}}}:\mathsf{Seq}(\mathcal{A})\to\mathbb{R}^{2}$ and ${\bm{\mathrm{g}}}:\mathcal{A}\to\mathbb{R}^{2}$, and $\mathcal{A}=\{y_{0},y_{1},y_{2}\}$. Let ${\bm{\mathrm{g}}}(y_{0})=(1,0)^{\top}$, ${\bm{\mathrm{g}}}(y_{1})=(1,1)^{\top}$, and ${\bm{\mathrm{g}}}(y_{2})=(1,-1)^{\top}$ be unembeddings, which do not fulfill the diversity condition (Definition 1): In fact, these unembeddings give $\mathcal{G}=\mathrm{span}({\bm{\mathrm{e}}}_{2})$, and $\dim(\mathcal{G})=1$ which is less than the dimensionality of the representation space. The vector ${\bm{\mathrm{e}}}_{2}:=(0,1)^{\top}$ is drawn as a blue arrow in Figure 4. We can construct another model where $\tilde{\bm{\mathrm{g}}}={\bm{\mathrm{g}}}$ and choose $\tilde{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})=({\bm{\mathrm{f}}}_{1}({\bm{\mathrm{x}}})+0.2\cos(40a_{1}/\pi),{\bm{\mathrm{f}}}_{2}({\bm{\mathrm{x}}}))^{\top}$. Figure 4 shows this transformation. By construction, this model generates the same next-token distribution of the first one; however, the two model representations are not equal up to a linear transformation.
Model $({\bm{\mathrm{f}}},{\bm{\mathrm{g}}})$ aba Model $(\tilde{\bm{\mathrm{f}}},\tilde{\bm{\mathrm{g}}})$
Refer to caption Refer to caption
Figure 4: Allowed distorsions among $\sim_{EL}$-equivalent models. From the left, model embeddings ${\bm{\mathrm{f}}}$ are given different colors. The red segment is non-linearly transformed on the right along $\tilde{\bm{\mathrm{f}}}_{1}$, whereas they remain equal to the left on the component $\tilde{\bm{\mathrm{f}}}_{2}$. This shows that the models are not $\sim_{L}$-equivalent.
B.6 Proof of Theorem 5
@@
For two matrices ${\bm{\mathrm{A}}}\in\mathbb{R}^{m\times k}$ and ${\bm{\mathrm{B}}}\in\mathbb{R}^{k\times n}$, from the rank-nullity theorem [Axler, 2015, page 62], it follows that:
@@
    Starting from log-equality of probabilities, we adopt a pivoting strategy to get rid of normalizing constants [^7] A similar “pivoting” strategy is the starting point of several identifiability proofs in nonlinear ICA with auxiliary variables, e.g., [Hyvarinen et al., 2019, Khemakhem et al., 2020b, a, Roeder et al., 2021].;
@@
Step 2: Obtaining the relation between embeddings ${\bm{\mathrm{f}}}$ and $\tilde{\bm{\mathrm{f}}}$. We will now show that ${\bm{\mathrm{M}}}:=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}({\bm{\mathrm{G}}}^{\top})^{+}\tilde{{\bm{\mathrm{G}}}}^{\top}({\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}})$ and $\tilde{{\bm{\mathrm{M}}}}:=({\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}})^{+}(\tilde{\bm{\mathrm{G}}}^{\top})^{+}{{\bm{\mathrm{G}}}}^{\top}{\bm{\mathrm{P}}}_{{\mathcal{G}}}{\bm{\mathrm{P}}}_{{\mathcal{F}}}$ satisfies Equation (9), showing that ${\bm{\mathrm{f}}}$ and $\tilde{{\bm{\mathrm{f}}}}$ are linearly related. For this, consider $q$ elements, or tokens, $y\in\mathcal{A}$ such that the matrix
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{G}}})}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{f}}}({\bm{\mathrm{x}}})$ $=({\bm{\mathrm{G}}}^{\top})^{+}\tilde{{\bm{\mathrm{G}}}}^{\top}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}\tilde{{\bm{\mathrm{f}}}}({\bm{\mathrm{x}}})$ (From Equation (32), $({\bm{\mathrm{G}}}^{\top})^{+}{\bm{\mathrm{G}}}^{\top}={\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{G}}})}$)
@@
    We show that $m\leq k$ from the definition of ${\bm{\mathrm{M}}}$ in Equation (102);
@@
    We show that necessarily $m\geq k$ and $k^{\prime}\geq k$ from Equation (101);
@@
    We show that $k^{\prime}\leq k$ from Equation (103).
(I) By equation (102), we have that ${\bm{\mathrm{M}}}={\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{M}}}$, since by Equation (102), in ${\bm{\mathrm{M}}}$ we have on the left ${\bm{\mathrm{P}}}_{\mathcal{M}}({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}=({\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{P}}}_{\mathcal{F}})^{+}$ by Lemma 2 (ii). Taking the rank of ${\bm{\mathrm{M}}}$ and using the fact that $\mathrm{rank}({\bm{\mathrm{A}}}{\bm{\mathrm{B}}})\leq\min\big{(}\mathrm{rank}({\bm{\mathrm{A}}}),\mathrm{rank}({\bm{\mathrm{B}}})\big{)}$ [Axler, 2015], we obtain:
@@
Together with 105, it shows that $k\leq m\leq k$, so it must be that $m=k$ and so $\mathrm{rank}({\bm{\mathrm{M}}})=\mathrm{dim}(\mathcal{M})$.
@@
which, together with 110, it holds only as long as $k^{\prime}\leq k$. This shows that $k\leq k^{\prime}\leq k$, so it must be that $k^{\prime}=k$.
@@
Step 4. Obtaining the relation between unembeddings ${\bm{\mathrm{g}}}$ and $\tilde{\bm{\mathrm{g}}}$. We will now show that ${\bm{\mathrm{N}}}:=({\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}})^{+}({\bm{\mathrm{M}}}^{\top})^{+}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}$, using the matrix ${\bm{\mathrm{M}}}$ in Equation (102) from Step 2 satisfies Equation (10).
Similar to Step 2, we take $\ell$ points, or sequences, ${\bm{\mathrm{x}}}_{i}\in\mathsf{Seq}(\mathcal{A})$ such that ${\bm{\mathrm{F}}}$ and $\tilde{{\bm{\mathrm{F}}}}$ in Equation (106) span $\mathcal{F}$ and $\tilde{\mathcal{F}}$, respectively. We then have:
@@
${\bm{\mathrm{F}}}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}}{\bm{\mathrm{g}}}_{0}(y)$ $=\tilde{{\bm{\mathrm{F}}}}^{\top}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (Using orthogonal projectors Equation (97))
@@
where in the last line we substituted the expression for ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{F}}}$ given by Equation (102). Thus, restarting from (114), and reworking the expression we get:
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}}{\bm{\mathrm{M}}}^{\top}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}{\bm{\mathrm{g}}}_{0}(y)$ $={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{G}}}\tilde{\bm{\mathrm{g}}}_{0}(y)$ (From Equation (32) we get $(\tilde{{\bm{\mathrm{F}}}}^{\top})^{+}(\tilde{{\bm{\mathrm{F}}}}^{\top})={\bm{\mathrm{P}}}_{\mathrm{Im}(\tilde{{\bm{\mathrm{F}}}})}={\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}$)
@@
where we used in the last line that ${\bm{\mathrm{P}}}_{\tilde{\mathcal{F}}}{\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}{\bm{\mathrm{M}}}^{\top}={\bm{\mathrm{P}}}_{\tilde{\mathcal{M}}}{\bm{\mathrm{M}}}^{\top}={\bm{\mathrm{M}}}^{\top}$, and ${\bm{\mathrm{P}}}_{\tilde{}\mathcal{M}}{\bm{\mathrm{M}}}^{\top}={\bm{\mathrm{M}}}^{\top}$ follows by the definition of ${\bm{\mathrm{M}}}$ (Equation (102)), containing on the right $({\bm{\mathrm{P}}}_{\tilde{}\mathcal{G}}{\bm{\mathrm{P}}}_{\tilde{}\mathcal{F}})$. Recall that by Lemma 2 (iii), it holds that
@@
Using this in Equation (115), we obtain
@@
where in the last line we use $({\bm{\mathrm{M}}}^{\top})^{+}{\bm{\mathrm{M}}}^{\top}={\bm{\mathrm{P}}}_{\mathrm{Im}({\bm{\mathrm{M}}})}={\bm{\mathrm{P}}}_{\mathcal{M}}$, where the first equality follows by Equation (32) and the second one is given by $\mathrm{Im}({\bm{\mathrm{M}}})=\mathcal{M}$ from Equation (113). We now use that ${\bm{\mathrm{P}}}_{\mathcal{M}}{\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}={\bm{\mathrm{P}}}_{\mathcal{F}}{\bm{\mathrm{P}}}_{\mathcal{G}}$ to obtain:
@@
    We use Equation (120) to obtain that $\mathrm{rank}({\bm{\mathrm{N}}})\geq k$.
@@
(II) Next, consider $q$ elements of $\mathcal{A}$ such that the matrices ${\bm{\mathrm{G}}}$ and $\tilde{{\bm{\mathrm{G}}}}$ in Equation (95) have rank equal to $\mathrm{dim}(G)$ and $\mathrm{dim}(\tilde{\mathcal{G}})$, respectively. From Equation (120), it holds: